NASA Logo - GSFC + NASA Homepage
+ Goddard Homepage

Known Issues with Gauge Data Rain Gauge: GPM Precipitation Science Research Facility

Known Issues with Gauge Data


April 2013

April 16-17, 2013
  • Disregard all tips between April 16-17.
  • Reset all Nassawadox gauge server ip's. Sent test packets. Cleaned solar panels, releveled gauges as needed, trimmed sites as needed and updated coordinates and elevations using Garmin model GPS 60 handheld device.
  • Removed NASA011 from Wallops PRF pad and installed at Southern Tip. Tested OK.
NASA025
  • NASA025 had 1/2" water standing in tub. Drained water out of tub and in process, the gps coaxial connector became dislodged from the connector on the black box. We were able to reinsert it without any issues. Gps is working OK. We will keep an eye on it for a while.
  • Needs bird spikes.
NASA013
  • Metal screws need to be replaced with nylon ones.
  • Needs bird spikes.
NASA021
  • Trimmed around system.
  • Needs bird spikes and nylon screws.
NASA008
  • Trimmed around system.
  • Needs bird spikes and NASA sign.
NASA015
  • Found that both gauge cans were slightly askew on top of the gauge bases. It appears that B may have blown off and someone found it and put it back. Gauge B is missing the bird spike and small screen in the can. Will need to replace. Trimmed around system. Reset server IP. Cleaned sp. Releveled gauges.
NASA007
  • Need to verify the A/B position of the two gauges.
NASA005
  • Cans were dirty. Cleaned and leveled gauges.
  • Needs bird spikes.
  • Trimmed around system.
NASA018 (17 April 2013)
  • Replaced solar charger in black box.
NASA018 (22 April 2013)
  • NASA0018 still losing voltage. Will need to bring system into lab and replace sp and harness.
NASA022
  • Trimmed around system, cleaned cans.
  • Needs bird spikes.
April 30, 2013
NASA018
  • Brought NASA018 into the lab for diagnostics. Changed the battery to previous one.
  • Adjusted input power connector to PCB. Both female pins were not mating to male pins on PCB properly.
  • Thoroughly cleaned and leveled system, performed tip tests and packet test and manually changed server ip

May 2013

NASA018
May 13, 2013 (0930-1100)
  • After testing NASA018 for one week, reported battery voltage was significantly lower than normal. Paul and I decided to switch the black box with NASA019's to isolate whether the issue lies in the solar panel or the circuitry. NASA019 will be located on the Wallops Island test site (37.93706, -75.47349) until troubleshooting has been resolved.
  • We will continue to monitor both systems for any changes.

June 2013

Disregard any and all gauge tips between June 4 - June 6 maintenance period. Disregard any and all tips between 1300z and 1430z on 12 June 2013.
NASA013
June 4, 2013 (1155-1230)
  • Both A and B gauge funnels were significantly clogged with bird debris. Cleaned and leveled gauges as needed.
  • Installed bird spikes on both gauges.
  • Replaced outer can stainless steel screws with black Polypropylene screws.
NASA021
June 4, 2013 (1315-1347)
  • Found site very dirty with black powdery fine dirt over everything.
  • Installed bird spikes on both gauges.
  • Replaced outer can stainless steel screws with black Polypropylene screws.
NASA002
June 5, 2013 (0910-0924)
  • Gauge A funnel was clogged with debris. Unclogged, cleaned and leveled gauges as needed.
NASA008
June 5, 2013 (0910-0924)
  • Forgot to install bird spikes at this site. Will do so on next visit.
NASA025
June 5, 2013 (1020-1035)
  • Installed bird spikes at this site.
NASA006
June 5, 2013 (1035-1055)
  • Heavy layer of yellow pollen dust in both buckets.
  • Replaced 2 stuck black screws in gauge B outer can.
NASA011
June 5, 2013 (1057-1106)
  • Will need to install bird spikes at this site next visit.
NASA015
June 5, 2013 (1154-1205)
  • Gauge B bird spike missing as well as small screen. Replaced both.
  • Gauge B outer can orifice clogged up.
NASA018
June 6, 2013 (1004-1029)
  • Installed bird spikes.
  • NASA018 is temporarily located at Dr. Stith Property while we evaluate NASA0019 at the main base.
NASA005
June 6, 2013 (1100-1121)
  • Site was very dirty. Will need to install bird spikes next visit.
  • Both Gauge A and B outer can orifices were clogged up.
NASA022
June 6, 2013 (1131-1145)
  • Installed bird spike on gauge A only. Will have to install bird spike on gauge B next visit.
NASA001
June 6, 2013 (1205-1215)
  • Overall, site was clean. However both gauge A and B outer can orifices were clogged with debris.
NASA017
June 6, 2013 (1230)
  • Tried twice to visit Water Treatment Plant site. Gate was locked with no one around.
NASA003
June 6, 2013
  • Tried twice to visit Salt Marsh Cradle site. Chain was across lane and my key would not work in it. Have contacted Dave Harris to try to resolve situation..
  • Maintenance Report
NASA008
June 12, 2013
  • Inspected Gladden Farm site at 1029 local. Found that Gauge A bucket assy had come out of pivot points. Readjusted bucket assy. No other problems found.
NASA0013
June 12, 2013
  • Inspected Dairy Farm site at 1110 hours local. Found that Gauge B bucket assy had come out of pivot points. Readjusted bucket assy. No other problems found.
  • Maintenance report
NASA018-19
June 20, 2013
  • Robert Tittle and I removed NASA019 from Wallops Island and took it back to its original location at the Dr. Stith Property. We exchanged the NASA018 black box at the Stith property with NASA019. We then took NASA018 to its original location at the Hermitage Farms Plant Nursery and reinstalled it there. Both sites were tested and found to be working fine.Disregard any tips from 0920 to 1045 local for both these sites on this day.

August 2013

All tips starting on 1 August 2013 from below mentioned gauge systems should be ignored until further notice.
August 1, 2013 (0800-1430)
  • Removed NASA015, 007, 023 and 019 from their respective properties. Paul and I set them up at Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed. All tips starting today should be ignored until further notice.
  • All gauges are scheduled to be moved from Nassawaddox to Wallops Island over the next few weeks.
August 2, 2013 (0800-1400)
  • Removed NASA022, 002, 018 and 005 from their respective properties. Paul and I set them up at Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed. All tips starting today should be ignored until further notice.
August 20, 2013 (0800-1430)
  • Removed NASA024, 001, 014 and 008 from their respective properties. Paul and I set them up at Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed. All tips starting today should be ignored until further notice.
August 26, 2013
  • NASA 001, 002, 005, 007, 008, 014, 015, 018, 019, 022, 023, 024 are level and clean. All data from these gauges are valid as of now. The onboard microSD cards have been backed up.
  • GPS coordinates have been updated for the above mentioned platforms and the server ip address has been updated for each platform as well.
August 28, 2013
Disregard all tips from below mentioned platforms on this day.
  • Placed NASA0014 at the new Sysco location (N 38.08171 W -075.59929, elev. +- 4m).
  • Gauges are not level. Cleaning was performed, packet was sent and tip test was done. We could not get a positive gps lock. Will return tomorrow to reboot the system and try to get a good lock and also will level the gauges.
  • Placed NASA0018 at the Chip Mill site (N 38.07449 W -075.58916, elev. +- 3m). Gauges are level, clean, gps is locked and packet was sent. Ground rod was not able to be installed due to platform sitting on 6+ inches of concrete. We will return tomorrow to install an extension for the grounding rod.
August 29, 2013
Disregard all tips from below mentioned platforms on this day.
  • Placed NASA0024 at the new Doc Johnson location (N38.05513 W075.57075 elev. +- 2m). Gauges are level, clean, gps is locked and packet was sent.
  • Placed NASA0002 at the new Mike Sigrist Farm location (N38.08469 W 075.58570 elev. +- 4m). Gauges are level, clean, gps is locked and packet was sent.
August 30, 2013
Disregard all tips from below mentioned platforms on this day.
  • (1110-1140) Installed NASA0007 at John Gerlach's site.(N38.09977 W 075.55283 elev. +- 2m). Gauge serial numbers were not verified.
  • (1230-1300) Installed NASA0001 at the new Lowes site (N 38.07776 W 075.55817 elev. +- 3m). Gauge A M9594 Gauge B M9581.

September 2013

September 3, 2013
Disregard all tips from below mentioned platforms on this day.
  • (1140-1215)Installed NASA0008 at the new Williams Street Well site (N38.06687 W 075.57955 elev. +- 7m). Gauge A M9590 Gauge B M9287.
  • Arrived at Sysco site at 1230. Removed NASA0014 and replaced it with NASA0015. Everything tested ok. Need to come back with string trimmer and trim around the platform. Finished at 1258 hours.
September 4, 2013
  • Installed fencing around NASA0001 (Lowes Site).
  • Trimmed grass around NASA0015 at Sysco Site.
September 5, 2013
Disregard all tips from below mentioned platforms on this day.
(1050-1200)
  • Installed NASA0019 at new Atlantic Tractor Site (N 38.08171, W 075.58957 elev. +-9m) without problems. All tests checked good and installed sign.
    • Need to install PVC pipes (x2) in the future.
    • As per Jeff (General Manager), we do not need to call in advance before performing maintenance.
    • Gauge A: M9290, Gauge B: M9585
    • This location is in a different on site location than the one we were going to use. Due to high traffic in the original location we found a place about 150 yards north of the original site.
(1210-1245)
  • Installed NASA0005 at temporary site at Paul Bashor's house (N 38.09393 W 075.56766 elev. +- 4m). Disconnected and set aside the three Eastern Shore gauges (S11A, B and C) and installed NASA0005 in their place. All tests checked good. No sign was installed. Gauge A: M9586, Gauge B: M9299.
September 10, 2013
Disregard all tips from below mentioned platforms on this day.
(0950-1100)
  • Using the city's utility boat, we installed NASA0023 at the new Pocomoke Fairgrounds Island site (N38.0640, W 075.57008, Elev. +-0m). Gauges are level, GPS got a lock and surrounding area was cleaned of debris. We installed the system on the highest part of the island which may be about 30' across.
    • Point of contact is Chris Marlett, who is Bill East's 2nd in charge.
    • Chris mentioned that they have never had water come over the island.
    • Gauge A: M9591, Gauge B: M9295.
    • Chris Marlett mentioned that it would not be feasible to bring the utility boat out every time we wanted to do maintenance. Will probably have to use hip waders from now on.
    • Chris also gave us a key for the Fish Plant Site and the William Street Well Site.
    • No sign is installed at this site.
(1110-1153)
  • Installed NASA0022 at the new Fish Plant Site (N 38.05583, W 075.56447, Elev. +- 8m). As promised, Chris Marlett had his people bush hog the area we wanted to use. We also used a chainsaw to cut down some saplings and high bushes in the area. Gauges are level, GPS got a lock and surrounding area is clear of debris.
    • Gauge A: M10754, Gauge B: M9289.
    • Note: Gauge B needs bird spikes installed.
September 13, 2013
Disregard all tips from below mentioned platforms on this day.
(1040-1215)
  • Installed NASA0014 at the new American Legion Site (N38.05583 W075.54845 Elev. +-7m). Manually set coordinates since the gps is non-functional in the unit. Ran tip tests and sent a packet successfully. Installed fencing around site and took pictures.
    • gauges: A=M9302, B=M9304
September 16, 2013
Disregard all tips from below mentioned platforms until further notice.
(1000-1400)
  • Removed NASA0021 and NASA0013 from their respective properties in Nassawadox and installed at the Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed due to required cleaning and maintenance. All tips starting today should be ignored until further notice.
    • NASA0021 gauge A was severely clogged with debris.
September 18, 2013
Disregard all tips from below mentioned platforms until further notice.
(1000-1400)
  • Removed NASA0009 and NASA0010 from their respective properties in Nassawadox and installed at the Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed due to required cleaning and maintenance. All tips starting today should be ignored until further notice.
    • Found a small bubble under NASA0010s solar panel. No voltage drop has been noted.
September 19, 2013
Disregard all tips from below mentioned platforms until further notice.
(0900-1200)
  • Removed NASA0017 and NASA0004 from their respective properties in Nassawadox and installed at the Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed due to required cleaning and maintenance. All tips starting today should be ignored until further notice.
    • NASA0017 needs bird spikes installed.
(1245-1500)
  • Cleaned, leveled and performed successful diagnostics on NASA0004, NASA0009, NASA0010, NASA0013, NASA0017 and NASA0021. All tips from these platforms starting this day are valid.
September 20, 2013
Disregard all tips from below mentioned platforms until further notice.
(1100-1500)
  • Removed NASA0016 and NASA0020 from their respective properties in Nassawadox and installed at the Wallops Island test pad. Units are not leveled, grounded and no soil moisture/temp probes are installed due to required cleaning and maintenance. All tips starting today should be ignored until further notice.
September 23, 2013
Disregard all tips from below mentioned platforms on this day.
(1415-1440)
  • Installed NASA0009 at Hayward's Lott site(Coords=N38.09264, W075.58957, Elev. 7m). Gauges are level, GPS got lock and server ip was changed to alias. Ran tip tests and sent a packet successfully.
    • Gauge A = M9298 @ .01" resolution.
    • Gauge B = M10770 @ .01" resolution.
September 24, 2013
Disregard all tips from below mentioned platforms until further notice.
  • Cleaned, leveled and performed successful diagnostics on NASA0016 and NASA0020. Server I.P.s were changed to alias.
  • Changed I.P. address to alias on NASA0004, NASA0010, NASA0013, NASA0017 and NASA0021.
September 26, 2013
Disregard all tips from below mentioned platforms until further notice.
(1435-1500)
  • Arrived at Somerset road site (N38.08926, W -75.56734, Elev. + 1m) with NASA0020. Paul and Jason took pictures, leveled both gauges and ran successful diagnostics.
    • This site needs a sign installed.
    • Gauge 'A': M9301, .01" resolution. Gauge 'B': M10761, .01" resolution.
(1405-1420)
  • Arrived at Crip Taylor site (N38.08543 W -75.57449 Elev. + 2m) with NASA0016. Took pictures, leveled both gauges and ran successful diagnostics.
    • Gauge 'A': M1300, .01" resolution. Gauge 'B': M10767, .01" resolution.
    • Need to speak with Crip to find out if it is ok to put up sign.
September 30, 2013
Disregard all tips from below mentioned platforms until further notice.
(1230-1315)
  • Arrived at James East Farm (N38.09748, W -75.5950, Elev. +13m) with NASA0021. Took pictures, leveled both gauges and ran successful diagnostics.
    • Gauge B tipping mechanism was stuck. Readjusted mechanism and it checked good.
    • Need to go back and install sign.
(1340-1420)
  • Arrived at American Legion site (NASA0014) to replace defective GPS unit with new one. Tested and got successful GPS lock. Sent test packet, cleaned solar panel and re-leveled gauges.

October 2013

October 15, 2013
Disregard all tips from below mentioned platforms during specified time on this day.
(1000-1330)
  • Lightly cleaned and re-leveled NASA0005, NASA0020, NASA0016, NASA0014, NASA0001 and downloaded data from flash drives as needed.
  • NASA0020's tipping mechanism of gauge A had come out of pivot points. Re-adjusted tipping mechanism to fix problem.
  • NASA0016, 0014 and 0001 had a slight bit of water in the tub ( <2"). None of the black boxes was affected by water. No other issues were found.
October 16, 2013
Disregard all tips from below mentioned platforms during the specified time on this day.
(0825-1455)
  • Lightly cleaned and re-leveled NASA0007, NASA0005, NASA0021, NASA0009, NASA0002, NASA0019, NASA0015, NASA0018, NASA0020, NASA0016, NASA0001 and NASA0014. Downloaded data from flash drives. Replaced bird spike wire with zip ties as needed.
  • NASA0021 and NASA0001 had a slight bit of water in the tub ( <2"). None of the black boxes was affected by the water. No other issues were found.
October 17, 2013
Disregard all tips from below mentioned platforms during the specified time on this day.
(1150-1427)
  • Lightly cleaned and re-leveled NASA0008, NASA0022 and NASA0024. Downloaded data from flash drives. Replaced bird spike wire with zip ties as needed.
  • NASA0024 gauge 'B' had come out of it's pivot points. Readjusted and checked good. No other issues were found.
October 18, 2013
Disregard all tips from the below mentioned platform during the specified time on this day.
(1010-1050)
  • Arrived at Pocomoke Fairgrounds site (NASA0023) at 1010 hours local. Downloaded data from flash drive, cleaned and leveled gauges as needed. Dave Marks accompanied us and took pictures for documentation. No issues found.
October 22, 2013
Disregard all tips from the below mentioned platforms during the specified times on this day.
(1250-1410)
  • Jason, Paul and Dave Wolff arrived at the Pocomoke Middle School (N 38.06717, W -075.56688, Elev. + 10m)at 1250 hours local. We installed NASA0004. Diagnostics were done and all tests checked good. Installed fence around gauge site. Installed signage. No other issues pending.
    • Gauge 'A' : M9297, resolution: .01".
    • Gauge 'B' : M9308, resolution: .01".
    • Need to go back and install protective t-post covers for fence.
(1430-1520)
  • Jason, Paul and Dave Wolff arrived at the Pocomoke High School (N 38.06023, W -075.55296, Elev. + 9m)at 1430 hours local. We installed NASA0013. Diagnostics were done and all tests checked good. Installed fence around gauge site. Installed signage. No other issues pending.
    • Gauge 'A' : M9583, resolution: .01".
    • Gauge 'B' : M10759, resolution: .01".
    • Need to go back and install protective t-post covers for fence.
October 28, 2013
Disregard all tips from below mentioned platforms during the specified times on this day.
(1030-1105)
  • Arrived at the Josh Lankford Farm (N 38.08731, W -075.59969, Elev. +4m) at 1030 hours local. Installed NASA0010. Diagnostic tests were performed and all tests were good. No other issues found.
    • Gauge 'A': M9296, resolution: .01".
    • Gauge 'B': M9584, resolution: .01".
    • The outer can label has the incorrect resolution of .2mm. The gauge itself was verified as being a .01" resolution by MetOne. Replacement label needs to be applied.
(1350-1520)
  • Arrived at the Pocomoke City Visitor Center (N 38.07595, W -075.57011, Elev: +4m) at 1350 hours local. Installed NASA0017 on the roof of the visitor center. Gauge is mounted sitting on a sheet of 1/2" plywood, and with nylon strapping fastened to roof seams for extra stability. Gauge was tested and all tests checked good. No sign was put on the roof.
    • Gauge 'A': M10773, resolution: .01".
    • Gauge 'B': M9292, resolution: .01".
October 29, 2013
Disregard all tips from below mentioned platforms during the specified times on this day.
(1420-1435)
  • Arrived at Chip Mill site (NASA0018) at 1420 hours local. Fastened a length of copper wire using ground lugs to the existing ground wire and secured it in the ground. Re-leveled gauges as needed.
(1447-1452)
Lowes site (NASA0001)
  • Installed orange t-post protectors on the fence of the gauge site there.

November 2013

November 1st, 4th and 5th 2013
Disregard all tips from below mentioned platforms until further notice.
(All Day)
  • Removed NASA0003 (Salt Marsh Cradle, Nov 1st), NASA0012 (Colonial House, Nov 4th), NASA0025 (North Solar Panel, Nov 4th), NASA0006 (South Solar Panel, Nov 5th) and NASA0011 (Southern Tip, Nov 5th) from their respective properties in Nasawaddox and staged them at the Wallops Island test pad for cleaning and diagnostics.
November 13, 2013
Disregard all tips from the below mentioned platform during the specified time on this day.
(1000-1035)
  • Dave Marks and Jason Bashor arrived at the James East farm (NASA0021) at 1000. Battery voltage had dropped significantly over the course of three days. Changed the battery and performed diagnostics. Diagnostics indicated that the new battery had a charge of 12.5 VDC. Leveled gauges and cleaned solar panel. No further issues.
November 15, 2013
Disregard all tips from the below mentioned platforms during the specified times on this day.
(1010-1027)
  • Arrived on site at Somerset Road Site (NASA0020). Removed old battery due to falling voltage and replaced it with a newer one. Ran diagnostics and everything tested ok. Releveled gauges as needed.
(1035-1048)
  • Arrived at Josh Lankford Farm (NASA0010). Removed old battery due to falling voltage and replaced it with a newer one. Ran diagnostics and everything tested ok. Releveled gauges as needed.

December 2013

December 2, 2013
Disregard all tips from the below mentioned platforms during the specified times on this day.
(1134-1153)
  • Arrived at Doc Johnson site (NASA0024) to replace the battery. We have been monitoring the battery voltage levels and noticed it has been dropping to as low as 11.4 VDC. The battery was replaced without any problems. Releveled the gauges, ran diagnostics and tip tests and everything tested ok.
(1338-1412)
  • Arrived at the Chip Mill Site (NASA0018) at 1338. The gauge was relocated to a location approximately 300 meters south-southeast of it's original position due to grid requirements. The updated coordinates are: [N 38.07211, W 075.58720, elevation: 15m.] Re-leveled gauges, reinstalled the the sign and ran succesful diagnostics.
(1422-1430)
  • Arrived at the Sysco site (NASA0015) to check the disparity between the tipping buckets. Found both gauges level and free tipping, with no blockages. Ran successful tip tests on both gauges. No reason for the disparity was found. We will monitor the gauge system closely in the future.
December 3, 2013
Disregard all tips from the below mentioned platforms during the specified times on this day.
(1103-1114)
  • Arrived at Hayward's Lott (NASA0009) to check the disparity between the tipping buckets. Found both gauges level and free tipping, with no blockages. Ran successful tip tests on both gauges. No reason for the disparity was found. We will monitor the gauge system closely in the future.
(1300-1500)
  • Removed inactive gauges NASA0003 and NASA0025 from the N159 pad for cleaning. Ran diagnostics on both gauges as well as tip tests. All tests were good. Replaced gauges on the N159 Pad at approximately 1500 hours local. Regard tips for these two gauges from 1500 to present.
December 4, 2013
Disregard all tips from the below mentioned platform during the specified time on this day.
(1043-1210)
  • Arrived at Harry Williams Site (N 38.07542, W 075.54588, elev: 4m) at 1043 hours local. Installed NASA0003. Diagnostics were done and all tests checked good.
    • This site needs a sign, bird spikes and a fence installed.
    • Gauge 'A' : MM10760, .01" resolution. Gauge 'B' : M10769, .01" resolution.
December 7, 2013
Disregard all tips from the below mentioned platform during the specified time on this day.
(1720-1750)
  • Dave Wolff noticed Gauge 'A' on NASA0009 at the Hayward's Lott Site was not tipping during the early event this day. Dave checked for blockages but did not see any. Performed a tip test several times to verify proper operation. System is operating as advertised.
December 13, 2013
Disregard all tips from the below mentioned platforms during the specified times on this day.
(0940-1140)
  • Arrived at the Glenn Holland Site (N 38.06242, W 075.58559, elev: -1m) at 1050 local. Installed NASA0006 and ran successful diagnostics and tip tests.
    • Gauge 'A': M9593 .01", Gauge 'B': M9596 .01"
(1300-1333)
  • Arrived at Fairgrounds Site (NASA0023) at 1300. Found both gauge bird spikes were flattened and both gauges screens were on the ground. Found gauge 'B' severely clogged with bird droppings. Cleaned both gauges and solar panel, straightened bird spikes, and performed successful diagnostics.
December 16, 2013
Disregard all tips from the below mentioned platforms during the specified times on this day.
(0954-1013)
  • Arrived at William Street Well Site (NASA0008) at 0954 local. Installed bird spikes and a new battery due to steadily falling voltages. Cleaned and leveled gauges as needed. Diagnostics and tip tests were not performed due to lack of equipment.
(1128-1151)
  • Arrived at the Fish Plant Site (NASA0022) at 1128 local. Found gauge 'A' partially obstructed from debris. Cleaned the gauge orifice, installed a bird spike on gauge 'B' and replaced the battery due to steadily falling voltage. Cleaned and releveled gauges as needed, adjusted gauge 'A' tipping mechanism since it was loose. Diagnostics and tip tests were not performed due to lack of equipment.
(1218-1228)
  • Arrived at the Chip Mill Site (NASA0018) at 1218 local. Replaced the battery due to steadily falling voltage. Cleaned and releveled gauges as needed. Diagnostics and tip tests were not performed due to lack of equipment.
(1240-1300)
  • Arrived at Josh Lankford Site (NASA0010) at 1240 local. Found a significant amount of standing water in the bottom of the tub. Emptied the water and dried the tub. Replaced the battery due to steadily falling voltage. Cleaned the solar panel and gauges and releveled gauges as needed. Replaced the incorrect outer can information label with correct one from the factory. Diagnostics and tip tests were not performed due to lack of equipment.
December 24, 2013
Disregard all tips from the below mentioned platform during the specified time on this day.
(1010-1035)
  • Arrived at Somerset Rd site (NASA0020) at 1010 local. Replaced the battery due to steadily falling voltage. The new battery has a charge of 12.7 vdc. Performed successful tip tests and packet. Cleaned and releveled gauges as needed.